# SELF-EXPLAIN: Teaching Large Language Models to Reason Complex Questions by Themselves

### v1

@@ 5.2 Results
How similar are self-explanations and human-crafted CoTs?
@@
The common standard to measure the quality of machine-generated samples is how similar they are to human-crafted ones (17; 30). The more similar, the better the quality is. However, despite the superior performance of self-explanation in prompting, we find that they are very different from human-crafted CoTs in terms of ROUGE-L score, terminology used and length. We use MedCQA as testbed which provides high-quality human-crafted explanations. For terminology comparison, we extract terms in both kinds of CoTs through scispaCy  [^2] https://allenai.github.io/scispacy/ and calculate F1 score between the two terms lists. Results are shown in Fig. 2. We find in terms of content (measured by ROUGE-L and term coverage), self-explanation differs from human-crafted CoTs greatly given the average similarity is around 15%. The length distribution of self-explanation is more centric, while the human-crafted CoTs have more varied lengths. We attach some example self-explanations in Appendix D.
6 Influence on Model Confidence
@@
Figure 4: Average confidence of model on each option for a balanced dataset where the number of questions that each option is the correct answer is equal. The correct options in exemplars are fixed to the same choice in each sub-figure for more rigorous comparison. The selection bias is calculated as $\frac{1}{4}{\sum_{i=1}^{4}|q_{i}-0.25|}$ where $q_{i}$ is model’s confidence for each option.
Models are more confident when prompted by self-explanation.
@@
For exemplars selected for generating self-explanation, we use Text-Davinci-003 to compute the average $P(y|x,\pi^{\text{self}},\theta)$ which reaches 99.96%, which may support our estimation of Eq. 3 in Sec. 4. In comparison, for human-crafted explanation of the same exemplars, average $P(y|x,\pi^{\text{human}},\theta)$ is lower, reaching 89.05%. This implies that given $(x,y)$, self-explanation may be the most related context to elicit $y$. For inference with self-explanation as demonstrations in ICL, log probabilities are computed for correct and wrong model outputs prompted with self-explanation and human-crafted one. Results are shown in Fig. 3. We can observe that models’ output log probabilities with self-explanation are much higher than with human-crafted explanation, indicating greater model’s confidence in its output. This suggests self-explanation can be more acceptable and effective for LLMs to elicit reasoning.
@@
We next consider two metrics measuring the influence of prompting with self-explanation and human-crafted CoT on LLM. (1) Calibration: the model’s output confidence should indicate the correctness of answers (e.g., wrong answers have lower confidence). Well-calibrated output confidence can assist human users to determine whether to trust model’s outputs or to look for alternatives in high-stakes applications. (2) Selection bias: the model should give equal chance of each option for a balanced multi-choice QA dataset. The experiment implementations are detailed in Appendix C. We find that (1) Self-explanation improves calibration. Fig. 3(a) showcases the model is generally more calibrated when using self-explanation. Further results in Fig. 3(b) show that output confidence by using human CoT is not indicative especially when the question is debatable (i.e., using self-explanation and human CoT generate different answers). Such results are consistent among different LLMs (see Fig. 5 in Appendix). (2) Self-explanation reduces bias. We control the correct option of in-context exemplars to be the same (e.g., all the correct options are A) for rigorous comparison. Results in Fig. 4 show that in all cases, using human-crafted CoTs leads to higher selection bias than using self-explanations.
7 Conclusion
In this work, we delve into whether LLMs can teach themselves to reason complex questions. Chain-of-Thought (CoT) can be costly to craft by humans and may have great uncertainty and noise. Therefore, we propose SELF-EXPLAIN inspired by encoding specificity hypothesis in memory retrieval to prompt LLMs to elicit reasonings given a question-answer pair. We find that when doing few-shot prompting with self-explanations as in-context CoT demonstrations, the LLM can surprisingly reach better performance than prompting with human-crafted CoTs on complex questions that require reasoning. We also find that LLM is more calibrated and less biased when prompting with self-explanations. However, comparison on four metrics shows that self-explanation is very different from human-crafted CoTs. Our work may thus challenge the common wisdom that human-crafted CoTs are the gold standard for prompting and possibly shed new light on understanding the effects of exemplars on ICL.
Appendix A Effects of Cues in SELF-EXPLAIN
In this section, we examine the performance of ICL with self-explanations prompted by different cues in our framework. We mainly follow cues in Liévin et al. [2022] as shown in Table 3. Since Liévin et al. [2022] focuses on medical domains, for general domains, we modify its cues by removing information specific to medical domains. We then generate self-explanations and perform ICL with them following our SELF-EXPLAIN framework. The final test results are shown in Table 4. We find no matter what cues are employed, ICL with self-generations elicited can all outperform using human-crafted CoTs, which demonstrates the robustness of SELF-EXPLAIN framework on the choice of cues.
Appendix B Effects of Generation Diversity
In this section, we examine the effects of generation diversity in SELF-EXPLAIN framework. By default, for each input of the exemplar, the LLM is prompted to generate five different ways of explanations to improve the generalization to test cases. At test time, one of five self-explanations is randomly chosen for an exemplar. To better understand the effects of this component in SELF-EXPLAIN, we therefore experiment with prompting different numbers of self-explanations for one exemplar input. Results are shown in Fig. 6. We find more diverse self-explanations to choose for an in-context exemplar can generally improve the test performance, while such improvement experiences diminishing return with further increased number of generations.
Appendix C Experiment Implementation Details for Sec. 6
We use MedCQA as testbed which offers human-crafted explanations for all the data. We use text-davinci-003 since it provides access to output probabilities. For the experiments on selection bias, we modify the correct options in in-context exemplars to make them the same in each case, e.g., all correct options are option A so that to give better control of influence of ICL. We append each option to the question to get the model’ confidence of the option token.
Appendix D Example self-explanation

### v3

@@ 4.2 Few-shot Prompting with Self-explanations
Prompting with self-explanations is better than using human-crafted CoTs.
Our results are shown in Table 1. Self-Explain can impressively outperform using CoTs crafted by human professionals by around 2% in both challenging MedQA and general domain, while reaching similar performance to Human CoT for MedCQA. Our approach also outperforms Auto-CoT (Zhang et al., 2023) and vanilla zero-shot CoT (Kojima et al., 2022), both of which cannot effectively surpass Human CoT for knowledge reasoning. The superior performance of Self-Explain may support our encoding specificity hypothesis. Overall, considering the difficulty and expense of crafting CoTs by humans, Self-Explain can thus be very useful in expert domains. Example self-explanations and human-crafted CoTs can be found in Appendix D.
Generation diversity is helpful.
Apart from naive Self-Explain, we generate five different self-explanations for each in-context exemplar and randomly select one for ICL at test time (see details in Section 3.1). As shown in results of “w/ Multi-Exp” in Table 1, this approach further boosts the performance of Self-Explain to significantly surpass Human CoT by around 4% in all datasets. To better understand the effects of this component, we experiment with generating different numbers of self-explanations for one exemplar input. Results are shown in Figure 3. We find generating different self-explanations for an in-context exemplar can generally improve the test performance, while such improvement experiences diminishing returns with further increased numbers of generations.
Does the correctness of self-explanations matter?
A natural question raised in Self-Explain is what if the self-generated explanations are wrong since the generation process is not supervised by humans. We look into this question by providing the LLM with random wrong answers to generate misleading explanations of the input question (i.e., we use $(x,y_{\text{wrong}})$ in Equation 3). Those self-explanations with wrong answers i.e., $(x,\pi^{\text{self}}_{\text{wrong}},y_{\text{wrong}}$) are then used for prompting as in-context exemplars. The results are shown in Table 2. We find that the performance of prompting with self-explanation seems insensitive to its correctness. This result suggests that a correctly labeled dataset may not be necessary for Self-Explain prompting. Similar results on text classification are observed that label space is more important for ICL than label correctness (Min et al., 2022). We similarly speculate that what carries more weight is how self-explanations are phrased, as they should match the context seen during encoding relevant information according to our encoding specificity hypothesis. We look deeper into this hypothesis in Section 4.3.
4.3 In-context Teaching via Explanations
We have demonstrated LLMs can teach themselves with Self-Explain for better knowledge reasoning. We further extend this to study whether self-explanations of one model can be used as in-context exemplars to teach another model through ICL. Teaching through supervised learning on teacher’s generated data has been widely investigated (Zhao et al., 2024; Ho et al., 2023; Hsieh et al., 2023), which can be framed as knowledge distillation. However, machine supervision through ICL has not yet been well studied. In this section, we have a teacher LLM generate self-explanations that are then used as in-context exemplars to teach a student LLM for reasoning unseen test cases. Saha et al. (2023) have explored a similar research question, while they insert teacher’s explanations into student’s generation for test examples during inference. This may not be fully considered as teaching as the taught model receives assistance with test examples, and its generalization ability is thus not evaluated.
The student is often its own best teacher.
Results are shown in Table 3. When doing few-shot prompting with the students’ own self-explanations as in-context exemplars, the students can generally reach the best performance, which is aligned with results in Section 4.2. This also supports the encoding specificity hypothesis. Noticeably, larger or stronger models may not necessarily be better in-context teachers. For example, for Llama2 and Mistral, prompting them with GPT-3.5’s explanations gives worse results than using those models’ own self-explanations. The performance for Llama2-7B with GPT-3.5 as the in-context teacher is even worse than not using any demonstrations of reasoning (i.e., “No CoT”).
Better teachers tend to produce rationales that more resemble student’s self-explanations.
The performance or scale of the teacher model is shown not indicative of its in-context teaching ability. Instead, we observe a strong correlation between students’ performance and the similarity between teachers’ demonstrations and students’ self-explanations. Specifically, ROUGE-L (Lin, 2004) is used to measure the closeness between rationales of teachers and students. We observe that explanations generated by different LLMs tend to be more similar to each other than to human-crafted rationales, as shown in Figure 4(a). Meanwhile, LLM teachers tend to yield better student’s performance than human teachers shown in Table 3. We further compute Pearson correlation coefficient between the student LLM’s test accuracy and the ROUGE score between its self-explanations and the teacher LLMs’ explanations/ human-crafted rationales. As shown in Figure 4(b), evident linear correlations between accuracy and ROUGE score are observed, especially for the models stronger than Llama2-7B. This may further support our encoding specificity hypothesis: in-context exemplars for the student should match its training data because rationales similar to the student’s self-explanations are more likely to align with the student’s training data.
Figure 5: Students’ accuracy improvement after applying Teach-Back. Values in brackets stand for students’ respective test accuracy w/ Teach-Back. The x-axis represents students who do reasoning. The y-axis is teacher models that provide in-context demonstrations for students.
4.4 Learning from Students for Better Teaching
Given the strong correlation between students’ test accuracy and closeness between teacher and student explanations, we are motivated to further look into the underlying causality, i.e., whether in-context teaching can be improved by letting the teacher learn the student’s self-explanations. In this section, the teacher model will first be fine-tuned on the student’s self-explanations to generate its new self-explanations that are then used as in-context exemplars for the student (i.e., Teach-Back introduced in Section 3.2).
Each student model generates self-explanations for 500 held-out training examples (will not be used for in-context demonstrations) for fine-tuning teacher models. To accommodate our available computing resources, we only fine-tune the teacher models whose size is smaller or equal to 7B with LoRA (Hu et al., 2021). Detailed implementations for fine-tuning are shown in Appendix B. Example generations before and after Teach-Back are shown in Appendix E.
Teach-Back improves in-context teaching.
As shown in Figure 5, when the teacher model is different from the student model, Teach-Back can greatly enhance the teaching performance of the teacher LLM, as evidenced by the improvement of test accuracy among students. Noticeably, fine-tuned teachers using Teach-Back can enable students to achieve significantly higher accuracy than the former best teachers in Table 3. For example, a fine-tuned Mistral-7B can guide Phi3-mini to achieve 59.3% accuracy. This is 4.9% higher than the accuracy achieved with an unfine-tuned Mistral-7B teacher and 1.6% higher than the best unfine-tuned teacher (i.e., GPT-3.5, see the column for “Phi3-mini” in Table 3). Interestingly, Teach-Back enables the smaller Mistral-7B to teach the much larger GPT-3.5 in context, surpassing human teachers by around 5% and Self-Explain by around 4% as visualized in Figure 7 of Appendix. Our results showcase the promising use of Teach-Back in leveraging a small tunable model to improve the few-shot prompting performance of a much larger LM without human supervision (i.e., human-crafted demonstrations).
Figure 6: Similarity between human-crafted CoTs and self-explanations in terms of ROUGE score, terminology covered, and length.
5 Further Analysis of LLMs’ Explanations
5.1 Faithfulness of Self-explanations
To elicit self-explanations $\pi^{\text{self}}$ from an LLM, we prompt it to explain a given pair of question and answer $(x,y)$ as shown in Equation 3, and then $(x,\pi^{\text{self}},y)$ will be used as one in-context demonstration. In this section, we evaluate how many of those raw self-explanations (before filtering) actually support the model to predict the given answer (Hase et al., 2020). We append the elicited explanations $\pi^{\text{self}}$ back to the given question $x$ as the prompt fed to the model. We then examine whether the model will correctly output the given answer $y$. We empirically find that diverse models can produce faithful explanations most of the time. For example, Mistral-7B reaches 94.2 % rate of faithful explanations, GPT 3.5 reaches 98.3% and Llama2-13B reaches 93.9 %. Our results may confirm the LLM’s ability to explain given questions and answers.
5.2 How Similar are Self-explanations to Human-crafted CoTs?
The common standard to measure the quality of machine-generated samples is how similar they are to human-crafted ones (Lu et al., 2022; Wang et al., 2022). The more similar, the better the quality is assumed to be. However, Hase et al. (2020) have pointed out that evaluation based on plausibility by matching human explanations is not sufficient. Our results also challenge this evaluation criterion. We show that LLMs’ self-explanations are very different from human-crafted CoTs in terms of ROUGE-L score, terminology used, and length. However, few-shot prompting with LLMs’ self-explanations demonstrates superior performance to using human-crafted CoTs.
@@ 5.2 How Similar are Self-explanations to Human-crafted CoTs?
We use MedCQA as our testbed, which provides high-quality human-crafted explanations. For terminology comparison, we extract terms in both kinds of CoTs through scispaCy  [^3] https://allenai.github.io/scispacy/ and calculate F1 score between the two terms lists. Results are shown in Fig. 6. We find in terms of content (measured by ROUGE-L and term coverage), self-explanation differs from human-crafted CoTs greatly, given the average similarity is around 15%. The length distribution of self-explanation is more centric, while the human-crafted CoTs have more varied lengths.
6 Related Work
In-context learning.
In-context learning is the ability of language models to induce answers from given demonstrations without weights updating in supervised tuning. In-context exemplars are the key to ICL which have dominating influence on the generation. Quite a few works have been proposed to optimize the selection of exemplars (Lu et al., 2023; Rubin et al., 2022; Fu et al., 2023b). On the other hand, in the cases of no access to task labels, Lyu et al. (2023) proposed a zero-shot ICL that employs pseudo exemplars with random labels for classification tasks.
Chain-of-Thought prompting without human-crafted exemplars.
Prompting with reasoning exemplars triggers LLMs to generate similar intermediate steps of thinking through ICL, known as Chain-of-Thought (CoT) (Wei et al., 2022). Kojima et al. (2022) propose zero-shot CoT prompting to elicit LLMs’ reasoning without human-crafted exemplars. This method is then leveraged to prompt LLMs to generate CoT exemplars by themselves for ICL (Zhang et al., 2023; Wan et al., 2023; Chen et al., 2023). Different from our work, which focuses on eliciting rationales from LLMs, Zhang et al. (2023); Wan et al. (2023); Shum et al. (2023) concentrate on selecting rationales generated according to Kojima et al. (2022). And Chen et al. (2023) further incorporate pseudo task generation alongside self-generated CoTs. Additionally, Yasunaga et al. (2024) propose analogical prompting to solve emerging new tasks without human-crafted demonstrations. Importantly, these works mainly focus on prompting engineering for very large, closed-source LMs (e.g., GPT-4). None of them formally investigate the fundamental in-context teaching among different LMs. Instead, our work proposes encoding specificity hypothesis to understand in-context teaching for LLMs, which is evidenced by experiments across different models.
Teaching via explanations.
Many past works have explored teaching student LLMs with teacher model’s explanations through supervised fine-tuning (Ho et al., 2023; Hsieh et al., 2023; Fu et al., 2023a). Few have investigated in-context teaching. Lampinen et al. (2022) demonstrate that LLMs can learn from human-crafted explanations in context. Instead of leveraging in-context exemplars, Saha et al. (2023); Lee et al. (2024) directly feed the teacher’s explanations of test examples to the student model during inference. In this case, the taught model receives direct assistance with test examples, and thus, its generalization ability from the teacher is not evaluated.
7 Conclusion
In this work, we investigate in-context teaching, where a teacher provides in-context example rationales to teach a student to reason over unseen questions. We introduce the encoding specificity hypothesis that in-context exemplars at test time should match the student model’s related training examples. Motivated by our hypothesis, we propose Self-Explain to let an LLM teach itself with its self-explanations through in-context learning, which outperforms human-crafted chain-of-thoughts and other baselines in different reasoning tasks. We reveal that for in-context teaching, rationales by distinct teacher LLMs or human teachers that more resemble the student LLM’s self-explanations are better demonstrations, which further supports the encoding specificity hypothesis. We then propose Teach-Back to align the teacher LLM with the student, which can enhance the in-context teaching performance.
8 Limitations
We propose Self-Explain and Teach-Back that verify our encoding specificity hypothesis for few-shot prompting. They also demonstrate impressive performance on diverse models for knowledge reasoning without human guidance. The student model’s performance with Self-Explain is consistently among the best. However, the student performance in Teach-Back does not necessarily surpass standard prompting with human CoTs, depending on the teacher model. For example, Mistral-7B with Teach-Back enables different student models to reach optimal test performance, while teachers like Llama2-7B are less effective. Therefore, we suggest using Self-Explain as a starting point in real applications. In the future, we will further investigate the influence of the teacher model in Teach-Back on student performance and how fine-tuning affects the teacher model’s self-explanations. Overall, in this work, the main contribution of our proposed Teach-Back is that it can greatly improve the ability of one LLM to teach a different student model.
In addition, our work is limited to only one teacher. Future work could explore many teachers, including mixture of experts. Moreover, there are various emerging advanced prompting methods for different kinds of reasoning tasks, e.g., tree-of-thoughts (Yao et al., 2024) or multi-round prompting (Khot et al., 2023; Zhou et al., 2022). In this work, we do not consider these more advanced designs of prompting, but focus on commonly used CoT prompting to eliminate the need of human-crafted CoTs. However, our approaches can be adapted to these methods e.g., by modifying the instructions used to elicit LLMs’ rationales. The majority of these methods still require human-crafted demonstrations. We will further investigate whether LLMs can implement these advanced prompting methods without human-crafted exemplars under our framework in the future.
Appendix A Effects of Instructions in Eliciting Self-explanations
In this section, we examine the performance of ICL with self-explanations prompted by different cues in our framework. We mainly follow cues in Liévin et al. (2022) as shown in Table 4. The first one is by default used in our framework. Since Liévin et al. (2022) focuses on medical domains, for general domains, we modify its cues by removing information specific to medical domains. We then generate self-explanations and perform ICL with them. The final test results are shown in Table 5. We find no matter what cues are employed, ICL with self-generations elicited can all outperform using human-crafted CoTs, which demonstrates the robustness of our proposed Self-Explain on the choice of cues.
Appendix B Implementation for Teach-Back with Fine-tuning
We reformat the training data with students’ self-explanations following the template in Table 6. We set the learning rate as $1\times 10^{-5}$ and fine-tune the teacher model with five epochs. We use the default setting for LoRA.
@@ Appendix B Implementation for Teach-Back with Fine-tuning
Appendix C Analysis on Model Confidence
@@
We find models are more confident with their predictions when using Self-Explain. For exemplars selected for generating self-explanation, we use Text-Davinci-003 to compute the average $P(y|x,\pi^{\text{self}},\theta)$ which reaches 99.96%. In comparison, for human-crafted explanation of the same exemplars, average $P(y|x,\pi^{\text{human}},\theta)$ is lower, reaching 89.05%. This implies that for a given $(x,y)$, the self-explanation $\pi^{\text{self}}$ provides a more relevant context for eliciting $y$ than the human explanation $\pi^{\text{human}}$.
@@
In addition, for inference with self-explanation as demonstrations in ICL, log probabilities are computed for correct and wrong model outputs prompted with self-explanation and human-crafted one. Results are shown in Figure 8. We can observe that models’ output log probabilities with self-explanation are much higher than with human-crafted explanation, indicating greater model’s confidence in its output. This suggests self-explanation can be more acceptable and effective for LLMs to elicit reasoning. Self-Explain may also improve model’s calibration. Calibration requires the model’s output confidence should indicate the correctness of answers (e.g., wrong answers have lower confidence). Well-calibrated output confidence can assist human users to determine whether to trust model’s outputs or to look for alternatives in high-stakes applications. Figure 8(a) showcases the model is generally more calibrated when using self-explanation. Further results in Fig. 8(b) show that output confidence by using human CoT is not indicative especially when the question is debatable (i.e., using self-explanation and human CoT generate different answers).
Appendix D Example self-explanations
@@ D.3 StrategyQA
Appendix E Examples of Different LLMs’ Self-explanations
Llama2-7B: The patient’s symptoms of decreased appetite, difficulty falling asleep, sadness, and compulsive behaviors (rechecking the door lock and counting toilet paper) are consistent with a diagnosis of obsessive-compulsive disorder (OCD). OCD is characterized by recurrent, intrusive thoughts (obsessions) and repetitive behaviors or mental acts (compulsions) that are time-consuming and cause significant distress or impairment.
The neurotransmitter serotonin is involved in the pathophysiology of OCD. Selective serotonin reuptake inhibitors (SSRIs), which increase the levels of serotonin in the synaptic cleft, are the most commonly used medications to treat OCD.
Mistral-7B: The patient’s symptoms are consistent with major depressive disorder and obsessive-compulsive disorder (OCD). OCD is characterized by recurrent and intrusive thoughts (obsessions) and repetitive behaviors or mental acts (compulsions) aimed at reducing anxiety or distress caused by the obsessions. Serotonin is a neurotransmitter that plays a crucial role in regulating mood and anxiety, as well as in the modulation of obsessive-compulsive symptoms. Therefore, pharmacotherapy for this patient should be targeted towards increasing serotonin levels in the brain. Dopamine is associated with motivation and reward, glutamate with learning and memory, and norepinephrine with the fight-or-flight response.
E.1 Examples of Self-explanations after Teach-Back
Mistral-7B fine-tuned on Llama2-7B: The patient’s symptoms are consistent with major depressive disorder and obsessive-compulsive disorder (OCD). Both conditions are associated with alterations in neurotransmission, particularly serotonin. Serotonin plays a role in mood regulation and impulse control, which are affected in major depressive disorder and OCD, respectively. Additionally, selective serotonin reuptake inhibitors (SSRIs) are commonly used to treat both conditions.
Llama2-7B fine-tuned on Mistral-7B: The patient’s symptoms are consistent with a diagnosis of obsessive-compulsive disorder (OCD). OCD is a neuropsychiatric disorder characterized by recurrent, intrusive thoughts (obsessions) and repetitive behaviors or mental acts (compulsions). The most commonly affected neurotransmitter in OCD is serotonin, which is involved in regulating mood, impulse control, and cognitive function. Therefore, pharmacotherapy targeting serotonin should be considered as a first-line treatment for the patient’s symptoms.
The other options are not as relevant to the patient’s symptoms:
A) Dopamine: While dopamine is involved in reward processing and motivation, it is not typically implicated in OCD.
B) Glutamate: Glutamate is involved in excitatory neurotransmission and is not typically associated with OCD.
C) Norepinephrine: While norepinephrine is involved in arousal and attention, it is not typically implicated in OCD.
In summary, the patient’s symptoms are best explained by a diagnosis of obsessive-compulsive disorder, and pharmacotherapy targeting serotonin is the most appropriate treatment option.